# Renormalisation for Reaction-Diffusion Systems with Non-Local Interactions

### v1

@@ 3.4 Critical Behaviour
where $X^{(0)}$ is the loop-less (tree) contribution to the density, $X^{(1)}$ is the single loop contribution, and so on. Note that the left hand side is asymptotic in $t$ with non-local interaction, as precision $\lambda<\infty$ is finite. On the right hand side, we have a small value $\gamma=\sqrt{\frac{t(\gamma)}{t}}$ to keep $t(\gamma)$ fixed and finite as $t$ gets large. From Eq. (14) this means that $g_{r}(\gamma)\rightarrow g^{*}$ approaches the fixed point, the precision $\lambda(\gamma)\rightarrow\infty$ gets large, suggesting all terms on the right hand side become identical to those seen in local interaction, returning the terms seen in [Lee1994], and implying the same universal behaviour is observed (to leading order, at least).
It is usually argued that above the critical dimension, the higher dimensionality means statistical fluctuations are negligible and only the uncorrelated residual mean field behaviour is relevant, where particles interact at the average rate of all others, as seen in tree level expansions. Although reasonable and intuitive, these descriptions are qualitative and when supra-critical effects are discussed in technical detail, they tend to revolve around limiting behaviour using renormalisation group ideas rather than statistical fluctuation negligibility [Binney1992, Hofmann2022].
@@
Now we have $g=R\kappa^{d-2}$ and scalings $g(\gamma)=g\gamma^{d-2}$, $n_{0}(\gamma)=n_{0}\gamma^{-d}$, meaning $Rn_{0}$ scales to $Rn_{0}\gamma^{-2}$. Thus as $\gamma\rightarrow 0$ these terms become large, and noting also that $\hat{R}(k,\lambda)\rightarrow 1+\mathcal{O}(\gamma^{2})$, under rescaling the $X^{(1)}$ term just reduces to the calculation in [Lee1994],
@@
Now for the case $\epsilon>0$ below the critical dimension, we find $g_{R}(\gamma)\rightarrow g^{*}$ as $\gamma\rightarrow 0$, and the same asymptotics as [Lee1994] result, with density decaying as $X(t)=\mathcal{O}(t^{-\frac{d}{2}})$.
@@ 4 Annihilating Walks with Birth and Death
The annihilation Model I in Section 3 only utilises some aspects of renormalization. Specifically, vertex functions can be summed in entirety and propagators require no normalisation (or more specifically, the vertex function $\Gamma^{(1,1)}$ is just a single diagram of the original propagator, no other diagrams start and finish with a single edge). As such it only uses additive renormalisation, no multiplicative renormalisation is needed. Next then, we consider full Model II (see Eq. (1)), which will requires additive and multiplicative methods of renormalisation. For this model, death, spontaneous birth, and branching are added to the annihilation process. Then any particle at position $p$ can spawn a new one at position $q$ in some non-local, isotropic fashion with rate Q(|pq|)Q(|p-q|). Additionally, a death rate $M_{p}$ and spontaneous creation rate $B_{p}$ is also supposed. This is a non local version of annihilating and branching random walks [Cardy1998]. The Liouvillian for Model II was given in Eq. (2), resulting in the action of Eq. (4).
@@ 4.1 Rescaling
Some comments are worth making. Firstly, the scaling factor $\eta$ only appears in the $D$ transform. This is in part due to the fact that all terms (such as $M_{p}=DM$, for example) have a $D$ factor (analogous to the choices made in [Janssen1999]). Again, like Model I, the natural choice to make is $\eta=\gamma^{2}$ which fixes diffusion, and renders time to have the same dimension as momentum squared. Secondly, the field renormalisation factors $\alpha$ and $\beta$ give some flexibility, although the quadratic and quartic terms are functions of $\alpha\beta=\gamma^{-d}$ and immune to any choice. Thirdly, notice that as $\gamma$ gets small, the variable $R^{(4)}$$^{\prime}$ will become small and irrelevant provided $d>2$. Thus it will not be a relevant variable near a critical dimension $d_{c}=4$, although, for low dimensions $d<2$, its effects would have to be considered. Following [Janssen2005, Tauber2005], it is natural to pick $\alpha$ and $\beta$ such that both cubic terms have the same coefficient and so scaling behaviour. Specifically, they are set as:
@@ 4.2 Tadpoles
The two main approaches to deal with the critical points for Model II with local interactions are (i) identify the critical point via divergence of the correlation function [Janssen2005, Tauber2014, Tauber2005] or (ii) implementation of an additive field transformation centered about the mean density $X_{p}=\braket{\psi_{p}}$ [Janssen1999]. Although the end product is the same, the second method is (subjectively speaking) perhaps a bit more transparent and so adopted for the non-local version of Model II. Thus the following field transformation is defined,
@@
Note, firstly, that for local interactions, this reduces to the expression given in [Janssen1999] once the irrelevant parameter $R$ is sent to zero (which will be seen to happen under rescaling). As we are considering steady state, the mean density $X_{p}=X$ is independent of space (as is the initial condition) and time, and so can be treated and written as a parameter.
@@
$G_{\bar{\phi}\phi}(k,\omega)$ $=\frac{1}{-i\omega+DF(k,\lambda)},$ $G_{\phi\phi}(k,\omega)$ $=\frac{2DgX\hat{Q}(k,\lambda)}{\omega^{2}+D^{2}F(k,\lambda)^{2}}.$
@@
Thirdly, we have the condition that the mean $\braket{\phi_{p}}=0$. Given that the $n_{0}$ term has been dropped, there are no initiating nodes at time $t=0$. Together, this tells us that the sum of all tadpole diagrams is zero. The first six diagrams are given in Fig. 6D. Note the tadpoles have no extruding arms such as the diagram given in Fig. 6C, which has an extra edge protruding to the right. If we replace this edge by all the tadpole terms, we have a sum of diagrams including this one. However, they must sum to zero as they have the sum of tadpoles (which is zero) as a factor. Thus any tadpoles that are not 1PI can be dropped (without having to define an effective action) [Peskin1995]. Note that as the particle density is uniform across space, the external momentum in diagrams is zero, meaning the only non-zero momenta present are circulating internal loops.
Now for local interaction, this results in the following (see [Janssen1999] for details),
@@
where now $\epsilon=d_{c}-d=4-d$, and following [Janssen1999], the parameter definitions below have been adopted ($I_{1}$ and $I_{2}$ are numerical constants [Janssen1999]),
@@
For the case of local interaction with $\hat{R}(k,\lambda)=\hat{Q}(k,\lambda)=1$ this reduces to that seen in [Janssen1999], meaning $F(k,\lambda)$ is quadratic and the primitive degree of divergence is $d=2$ as expected.
@@
For the case of local interaction, note from Eq. (20) that the order $\mathcal{\bar{\tau}^{-\epsilon}}$ term implies divergence near the critical point $\bar{\tau}\rightarrow 0$ (for $\epsilon>0$). This emerges as a result of the last two diagrams in Fig. 6D (see appendix of [Janssen1999] for details). This phenomenon can also be seen to occur for non-local interactions. For example, consider the penultimate diagram in 6D, which evaluates to the following ($\lambda$ dependencies such as $\hat{R}(k)\equiv\hat{R}(k,\lambda)$ are not shown to ease notation),
$I$ $=-2(Dg)^{3}\int_{0<t_{2}<t_{1}<t}\fint_{k}\fint_{l}\hat{R}(k)\hat{R}(l)^{2}G_{\bar{\phi}\phi}(k,t-t_{2})^{2}G_{\bar{\phi}\phi}(k,t_{2}-t_{1})G_{\phi\phi}(k-l,t_{2}-t_{1})G_{\phi\phi}(l,t_{2}-t_{1}).$
@@ 4.3 Renormalisation
The approach taken then, is as follows. The local interaction renormalisation scheme of [Janssen1999] is adopted. We then use the rescaling methods of Section 3 to approach the critical point, where again we find that solutions to Callan-Symanzik equations can be found directly. Then as we scale towards the critical point, we also find that non-local interactions become local, renormalisation is achieved and the same universality is obtained.
The renormalisation scheme is thus given as follows. We use a subscript 0 to denote the unrenormalised parameters (that is, the unrenormalised parameters above used thus far). Variables without subscripts are renormalised variables. This is taken from [Janssen1999] and included for convenience;
@@
Recall that because the mean density $X$ can be viewed as a parameter, independent of space-time in steady state, it can accordingly be renormalised. Note also that the last equation is replacing $g$ with a dimensionless parameter $u$, where $\mu$ is an arbitrary momentum scale, and the term $G_{\epsilon}=\frac{\Gamma(1+\frac{\epsilon}{2})}{(4\pi)^{\frac{d}{2}}}$, where $\epsilon=d_{c}-d=4-d$. The individual factors to order $\mathcal{O}(u^{2})$ are given by (again from [Janssen1999], although note that $g\rightarrow\frac{g}{2}$ and $u\rightarrow\frac{u}{4}$ are needed to get those results, as the actions are defined differently),
@@
$g(\gamma)=\gamma^{-\frac{\epsilon}{2}}g_{0}Z(u(\gamma))^{\frac{1}{2}}Z_{D}(u(\gamma))Z_{g}(u(\gamma))^{-1}=\gamma^{-\frac{\epsilon}{2}}g_{0}\left(1-\frac{3u(\gamma)}{\epsilon}+\mathcal{O}(u(\gamma)^{2})\right),$ (24)
@@
Now we see that (for $\epsilon>0$) as $\gamma\rightarrow 0$, $u(\gamma)\rightarrow\frac{\epsilon}{6}$, and as $\gamma\rightarrow\infty$, $u(\gamma)\rightarrow 0$. These limits flip for $\epsilon<0$. These are the fixed points reported in [Janssen1999]. Either way, with $\gamma=1$ we recover the original parameter $u$.
@@
which gives an implicit equation for the variation of $b(\gamma)$. As we approach the fixed point $\gamma\rightarrow 0$, the solution to the Callan-Symanzik equation given in [Janssen1999] can be found. More specifically, from Eq. (27) we find,
@@
$=$ $b(\gamma)\left(-\frac{d}{2}-2-\frac{u(\gamma)}{4\epsilon}\frac{\partial u(\gamma)}{\partial\log\gamma}\left(\frac{7}{4}-\frac{17}{2}\log\frac{4}{3}\right)\right),$
@@
Implicit in this equation is the value of $\mu$. This just sets the value of the unscaled renormalised parameters (that is, where $\gamma=1$). Now, although $\gamma$ is dimensionless, $\mu$ has the dimensions of momentum. If we use $\mu$ to make the arguments on the right hand side of Eq. (28) dimensionless [Barenblatt1996], we get an expression analogous to Eq. (19) in [Janssen1999], where, for small $\gamma$,
@@
Note, however, that this expression also includes a precision parameter and its scaling. Finally, then, we can consider the behaviour near the critical point. Now from the rescaling in Eq.s (28, 29) we note that as we approach the critical point with $\gamma\rightarrow 0$ the precision diverges and like Model I we find that the non-local interactions become local. At this point the derivation is identical to that in [Janssen1999], and the same universal behaviour is observed (see [Janssen1999] for details). Note that we have not derived or used the Callan-Symanzik to do this.
@@ 5 Conclusions
The scaling approach we have adopted can also give some context to the choice of rescaling parameters (such as $\kappa$ and $\mu$), which can seem very abstract when first encountered in reaction diffusion systems. Although in [Lee1994] $\kappa$ is defined via a vertex function, in [Janssen1999] $\mu$ it is an arbitrary scaling. Such scaling parameters are often related to measurements in quantum field theory and therefore tend to have a more concrete meaning and can be somewhat easier to comprehend. Here the rescaling of parameters via $\gamma$ with a preserved action gives context to both models in [Janssen1999, Lee1994].
@@
Finally we note that the case of two interacting particles has been studied here. For many benchmark studies [Cardy1998, Lee1994, Tauber2014], the case of $k$ interacting particles is considered. It would be natural to attempt an extension this work to those cases, although an interaction function such as $R(r_{1},r_{2},\dots,r_{k})$ would likely be a somewhat complicated exercise, even for spatially homogeneous systems.

### v2

@@ 3.4 Critical Behaviour
where $X^{(0)}$ is the loop-less (tree) contribution to the density, $X^{(1)}$ is the single loop contribution, and so on. Note that the left hand side is asymptotic in $t$ with non-local interaction, as precision $\lambda<\infty$ is finite. On the right hand side, we have a small value $\gamma=\sqrt{\frac{t(\gamma)}{t}}$ to keep $t(\gamma)$ fixed and finite as $t$ gets large. From Eq. (14) this means that $g_{r}(\gamma)\rightarrow g^{*}$ approaches the fixed point, the precision $\lambda(\gamma)\rightarrow\infty$ gets large, suggesting all terms on the right hand side become identical to those seen in local interaction, returning the terms seen in [23], and implying the same universal behaviour is observed (to leading order, at least).
It is usually argued that above the critical dimension, the higher dimensionality means statistical fluctuations are negligible and only the uncorrelated residual mean field behaviour is relevant, where particles interact at the average rate of all others, as seen in tree level expansions. Although reasonable and intuitive, these descriptions are qualitative and when supra-critical effects are discussed in technical detail, they tend to revolve around limiting behaviour using renormalisation group ideas rather than statistical fluctuation negligibility [4, 17].
@@
Now we have $g=R\kappa^{d-2}$ and scalings $g(\gamma)=g\gamma^{d-2}$, $n_{0}(\gamma)=n_{0}\gamma^{-d}$, meaning $Rn_{0}$ scales to $Rn_{0}\gamma^{-2}$. Thus as $\gamma\rightarrow 0$ these terms become large, and noting also that $\hat{R}(k,\lambda)\rightarrow 1+\mathcal{O}(\gamma^{2})$, under rescaling the $X^{(1)}$ term just reduces to the calculation in [23],
@@
Now for the case $\epsilon>0$ below the critical dimension, we find $g_{R}(\gamma)\rightarrow g^{*}$ as $\gamma\rightarrow 0$, and the same asymptotics as [23] result, with density decaying as $X(t)=\mathcal{O}(t^{-\frac{d}{2}})$.
@@ 4 Annihilating Walks with Birth and Death
The annihilation Model I in Section 3 only utilises some aspects of renormalization. Specifically, vertex functions can be summed in entirety and propagators require no normalisation (or more specifically, the vertex function $\Gamma^{(1,1)}$ is just a single diagram of the original propagator, no other diagrams start and finish with a single edge). As such it only uses additive renormalisation, no multiplicative renormalisation is needed. Next then, we consider full Model II (see Eq. (1)), which will requires additive and multiplicative methods of renormalisation. For this model, death, spontaneous birth, and branching are added to the annihilation process. Then any particle at position $p$ can spawn a new one at position $q$ in some non-local, isotropic fashion with rate Q(|pq|)Q(|p-q|). Additionally, a death rate $M_{p}$ and spontaneous creation rate $B_{p}$ is also supposed. This is a non local version of annihilating and branching random walks [5]. The Liouvillian for Model II was given in Eq. (2), resulting in the action of Eq. (4).
@@ 4.1 Rescaling
Some comments are worth making. Firstly, the scaling factor $\eta$ only appears in the $D$ transform. This is in part due to the fact that all terms (such as $M_{p}=DM$, for example) have a $D$ factor (analogous to the choices made in [19]). Again, like Model I, the natural choice to make is $\eta=\gamma^{2}$ which fixes diffusion, and renders time to have the same dimension as momentum squared. Secondly, the field renormalisation factors $\alpha$ and $\beta$ give some flexibility, although the quadratic and quartic terms are functions of $\alpha\beta=\gamma^{-d}$ and immune to any choice. Thirdly, notice that as $\gamma$ gets small, the variable $R^{(4)}$$^{\prime}$ will become small and irrelevant provided $d>2$. Thus it will not be a relevant variable near a critical dimension $d_{c}=4$, although, for low dimensions $d<2$, its effects would have to be considered. Following [18, 31], it is natural to pick $\alpha$ and $\beta$ such that both cubic terms have the same coefficient and so scaling behaviour. Specifically, they are set as:
@@ 4.2 Tadpoles
The two main approaches to deal with the critical points for Model II with local interactions are (i) identify the critical point via divergence of the correlation function [18, 32, 31] or (ii) implementation of an additive field transformation centered about the mean density $X_{p}=\braket{\psi_{p}}$ [19]. Although the end product is the same, the second method is (subjectively speaking) perhaps a bit more transparent and so adopted for the non-local version of Model II. Thus the following field transformation is defined,
@@
Note, firstly, that for local interactions, this reduces to the expression given in [19] once the irrelevant parameter $R$ is sent to zero (which will be seen to happen under rescaling). As we are considering steady state, the mean density $X_{p}=X$ is independent of space (as is the initial condition) and time, and so can be treated and written as a parameter.
@@
$G_{\bar{\phi}\phi}(k,\omega)$ $=\frac{1}{-i\omega+DF(k,\lambda)},$ $G_{\phi\phi}(k,\omega)$ $=\frac{2DgX\hat{Q}(k,\lambda)}{\omega^{2}+D^{2}F(k,\lambda)^{2}},$
@@
Thirdly, we have the condition that the mean $\braket{\phi_{p}}=0$. Given that the $n_{0}$ term has been dropped, there are no initiating nodes at time $t=0$. Together, this tells us that the sum of all tadpole diagrams is zero. The first six diagrams are given in Fig. 6D. Note the tadpoles have no extruding arms such as the diagram given in Fig. 6C, which has an extra edge protruding to the right. If we replace this edge by all the tadpole terms, we have a sum of diagrams including this one. However, they must sum to zero as they have the sum of tadpoles (which is zero) as a factor. Thus any tadpoles that are not 1PI can be dropped (without having to define an effective action) [27]. Note that as the particle density is uniform across space, the external momentum in diagrams is zero, meaning the only non-zero momenta present are circulating internal loops.
Now for local interaction, this results in the following (see [19] for details),
@@
where now $\epsilon=d_{c}-d=4-d$, and following [19], the parameter definitions below have been adopted ($I_{1}$ and $I_{2}$ are numerical constants [19]),
@@
For the case of local interaction with $\hat{R}(k,\lambda)=\hat{Q}(k,\lambda)=1$ this reduces to that seen in [19], meaning $F(k,\lambda)$ is quadratic and the primitive degree of divergence is $d=2$ as expected.
@@
For the case of local interaction, note from Eq. (20) that the order $\mathcal{\bar{\tau}^{-\epsilon}}$ term implies divergence near the critical point $\bar{\tau}\rightarrow 0$ (for $\epsilon>0$). This emerges as a result of the last two diagrams in Fig. 6D (see appendix of [19] for details). This phenomenon can also be seen to occur for non-local interactions. For example, consider the penultimate diagram in 6D, which evaluates to the following ($\lambda$ dependencies such as $\hat{R}(k)\equiv\hat{R}(k,\lambda)$ are not shown to ease notation),
$I$ $=-2(Dg)^{3}\int_{0<t_{2}<t_{1}<t}\fint_{k}\fint_{l}\hat{R}(k)\hat{R}(l)^{2}G_{\bar{\phi}\phi}(k,t-t_{2})^{2}G_{\bar{\phi}\phi}(k,t_{2}-t_{1})G_{\phi\phi}(k-l,t_{2}-t_{1})G_{\phi\phi}(l,t_{2}-t_{1})$
@@ 4.3 Renormalisation
The approach taken then, is as follows. The local interaction renormalisation scheme of [19] is adopted. We then use the rescaling methods of Section 3 to approach the critical point, where again we find that solutions to Callan-Symanzik equations can be found directly. Then as we scale towards the critical point, we also find that non-local interactions become local, renormalisation is achieved and the same universality is obtained.
The renormalisation scheme is thus given as follows. We use a subscript 0 to denote the unrenormalised parameters (that is, the unrenormalised parameters above used thus far). Variables without subscripts are renormalised variables. This is taken from [19] and included for convenience;
@@
Recall that because the mean density $X$ can be viewed as a parameter, independent of space-time in steady state, it can accordingly be renormalised. Note also that the last equation is replacing $g$ with a dimensionless parameter $u$, where $\mu$ is an arbitrary momentum scale, and the term $G_{\epsilon}=\frac{\Gamma(1+\frac{\epsilon}{2})}{(4\pi)^{\frac{d}{2}}}$, where $\epsilon=d_{c}-d=4-d$. The individual factors to order $\mathcal{O}(u^{2})$ are given by (again from [19], although note that $g\rightarrow\frac{g}{2}$ and $u\rightarrow\frac{u}{4}$ are needed to get those results, as the actions are defined differently),
@@
$g(\gamma)=\gamma^{-\frac{\epsilon}{2}}g_{0}Z(u(\gamma))^{\frac{1}{2}}Z_{D}(u(\gamma))Z_{g}(u(\gamma))^{-1}=\gamma^{-\frac{\epsilon}{2}}g_{0}\left(1-\frac{3u(\gamma)}{\epsilon}+\mathcal{O}(u(\gamma)^{2})\right).$ (24)
@@
Now we see that (for $\epsilon>0$) as $\gamma\rightarrow 0$, $u(\gamma)\rightarrow\frac{\epsilon}{6}$, and as $\gamma\rightarrow\infty$, $u(\gamma)\rightarrow 0$. These limits flip for $\epsilon<0$. These are the fixed points reported in [19]. Either way, with $\gamma=1$ we recover the original parameter $u$.
@@
which gives an implicit equation for the variation of $b(\gamma)$. As we approach the fixed point $\gamma\rightarrow 0$, the solution to the Callan-Symanzik equation given in [19] can be found. More specifically, from Eq. (27) we find,
@@
$=$ $b(\gamma)\left(-\frac{d}{2}-2-\frac{u(\gamma)}{4\epsilon}\frac{\partial u(\gamma)}{\partial\log\gamma}\left(\frac{7}{4}-\frac{17}{2}\log\frac{4}{3}\right)\right)$
@@
Implicit in this equation is the value of $\mu$. This just sets the value of the unscaled renormalised parameters (that is, where $\gamma=1$). Now, although $\gamma$ is dimensionless, $\mu$ has the dimensions of momentum. If we use $\mu$ to make the arguments on the right hand side of Eq. (28) dimensionless [2], we get an expression analogous to Eq. (19) in [19], where, for small $\gamma$,
@@
Note, however, that this expression also includes a precision parameter and its scaling. Finally, then, we can consider the behaviour near the critical point. Now from the rescaling in Eq.s (28, 29) we note that as we approach the critical point with $\gamma\rightarrow 0$ the precision diverges and like Model I we find that the non-local interactions become local. At this point the derivation is identical to that in [19], and the same universal behaviour is observed (see [19] for details). Note that we have not derived or used the Callan-Symanzik to do this.
@@ 5 Conclusions
The scaling approach we have adopted can also give some context to the choice of rescaling parameters (such as $\kappa$ and $\mu$), which can seem very abstract when first encountered in reaction diffusion systems. Although in [23] $\kappa$ is defined via a vertex function, in [19] $\mu$ it is an arbitrary scaling. Such scaling parameters are often related to measurements in quantum field theory and therefore tend to have a more concrete meaning and can be somewhat easier to comprehend. Here the rescaling of parameters via $\gamma$ with a preserved action gives context to both models in [19, 23].
@@
Finally we note that the case of two interacting particles has been studied here. For many benchmark studies [5, 23, 32], the case of $k$ interacting particles is considered. It would be natural to attempt an extension this work to those cases, although an interaction function such as $R(r_{1},r_{2},\dots,r_{k})$ would likely be a somewhat complicated exercise, even for spatially homogeneous systems.